Can someone help me please!!! Determine whether the equation defines y as a function of x
x^3 y-2y=6

Answer by stanbon(75887)

x^3 y-2y=6
---------
Yes, each value of x will have one corresponding value for y.
-------------------
y(x^3-2) = 6
----------------
At this point each value of x has one corresponding value of y
so y still a function of x.
y = 6/(x^3-2)
----------------
Comment: At this point the equation does not have
the same domain as the original equation.
================
Cheers,
Stan H.